OpenStudy (anonymous):

Find The Value Of: The Square Root Of The Quantity 18 Over 64.

OpenStudy (anonymous):

\[\sqrt{18/64}=\sqrt{9/32}=\sqrt{9/2*16}=\sqrt{3^{2}/2*4^{2}}=3/4\sqrt{2}=3\sqrt{2}/8\]

OpenStudy (anonymous):

THANK YOU!!(:

OpenStudy (anonymous):

yw